# ANote[fnsymbol]

### v2

@@ 5 Conclusion and Future Work
In this work, we propose an omni-perceive and -interactive tri-modal model, Nexus-O, that supports any combination of audio, image/video, and text inputs, and the Nexus-O-audio testbed. Unlike existing approaches that treat audio as an auxiliary, our model enables a joint understanding of these modalities. To facilitate the rapid deployment of the model in real-world scenarios, we investigate the synthesis of high-quality speech dialogue data augmented with actual environmental noise derived from public datasets. This approach substantially improves the performance and robustness of MLLMs in practical applications. To better understand the characteristics of diverse modalities throughout the integration process, we conducted an in-depth analysis from the perspective of representation space. As a cutting-edge contribution, we aim to establish a new standard for open-source multimodal foundation models and drive innovation in complex real-world applications.
In the future, we aim to enhance the capability of MLLMs to handle audio in a streaming manner, enabling real-time processing and generation of audio inputs and outputs for dynamic, time-sensitive applications. Additionally, we intend to integrate advanced Artificial Intelligence-Generated Content (AIGC) capabilities to allow the model to seamlessly generate high-quality, contextually coherent multimodal outputs, further broadening its applicability in creative and industrial domains.
@@ Appendix A Experimental Details
Appendix B Representational Alignment
Preliminaries.
Kernels, characterising the distance metrics between points in a representation space, are commonly used to assess vector space (Huh et al., 2024). Typically, the similarity between two kernels derived from different spaces indicates a higher degree of alignment between those spaces. This similarity can be quantified via kernel-alignment metrics, such as Centered Kernel Distance (cka) (Kornblith et al., 2019). Next, we formally define the target of our evaluation.
Given a data $x^{i}\in\mathcal{X}$ where $i=(1,\dots,N)$, each data can be represented as a corresponding tri-modal tuple $(x^{i}_{\text{language}},x^{i}_{\text{vision}},x^{i}_{\text{audio}})\in(\mathcal{X}_{\text{language}},\mathcal{X}_{\text{vision}},\mathcal{X}_{\text{audio}})$. For each data $x^{i}$, its vector representation can be mapped via feature function $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$. At the unembedding stage, $f$ for different modalities are the corresponding encoders (i.e., $f_{\text{language}}$, $f_{\text{vision}}$, and $f_{\text{audio}}$). At the embedding stage, $f$ is the unified LLMs with different modal inputs, denoted by $f_{\text{LLM}}$.
For each modality, a kernel, which measures the distance/similarity between datapoints, can be denoted by $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. The measurement of a single modal space can be described as:
$K(\mathcal{X})=K(x^{i},x^{j})=\langle f(x^{i}),f(x^{j})\rangle$
where $\langle$,$\rangle$ , ⟩ denotes the inner product operation, $x^{i}$ and $x^{j}$ are different input points. In this experiment, we specifically assess the alignment between language-vision modalities by measuring the alignment between kernels through kernel-alignment metrics, $m$:
$m(K(\mathcal{X}_{\text{language}}),K(\mathcal{X}_{\text{vision}}))$
To further evaluate whether audio modality can assist vision-language representational alignment at each stage, we fuse vision and audio spaces and examine whether this fusion leads to better alignment to language space. In detail, for the unembedding stage, we concatenate $f_{\text{vision}}(x^{i}_{\text{vision}})$ and $f_{\text{audio}}(x^{i}_{\text{audio}})$ as the fused feature space. The kernel can be described as:
$K(\mathcal{X}_{\text{vision}+\text{audio}})=\langle[f(x_{\text{vision}}^{i});f(x_{\text{audio}}^{i})],[f(x_{\text{vision}}^{j});f(x_{\text{audio}}^{j})]\rangle$ (1)
For the embedding stage, we feed both vision and audio inputs into the model. The kernel can be described as:
$K(\mathcal{X}_{\text{vision}+\text{audio}})=\langle f_{\text{LLM}}(x^{i}_{\text{vision}},x^{i}_{\text{audio}}),f_{\text{LLM}}(x^{j}_{\text{vision}},x^{j}_{\text{audio}})\rangle$ (2)
The final kernel-alignment scores can be described as:
$m(K(\mathcal{X}_{\text{language}}),K(\mathcal{X}_{\text{vision}+\text{audio}}))$
Figure 8: Tri-modal alignment visualisation at different hidden layers (blue: language, orange: vision, green: audio).
Figure 9: Tri-modal alignment visualisation at different hidden layers (blue: language, green: vision+audio).

### v3

@@ 5 Conclusion and Future Work
In this work, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline to overcome the current challenges, including limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. To this end, we propose an industry-level omni-perceive and -interactive tri-modal LLM, Nexus.
Extensive experiments validate the usefulness of our proposed pipeline and Nexus. In the visual understanding task, Nexus exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), Nexus is comparable to its backbone vocoder on Seed-TTS benchmark. In our real-world ASR testset, Nexus achieves outstanding performance, indicating its robustness in real scenarios.
In the future, we intend to link Nexus with vision generative models under our modularised framework to allow the model to seamlessly generate high-quality, contextually coherent visual outputs. Integrating the Artificial Intelligence-Generated Content (AIGC) capabilities into the current model can further broaden its applicability in creative and industrial domains, such as Embodied Intelligence.